OpenStudy (anonymous):

The number of computer shutdowns during any month has a poisson distribution, averaging 0.25 shutdowns per month. What is the probability of at least 3 computer shutdowns during the next year?

OpenStudy (perl):

.25 shutdowns per month = 3 shutdowns per year. you can use 3 as your lambda

OpenStudy (anonymous):

lamba is 3 or 3/12?

OpenStudy (perl):

lambda is 3 $$\Large { P(X\ge 3 ) = 1 - P( X < 3 ) \\= 1-(P(0)+P(1)+P(2)) }$$

OpenStudy (anonymous):

is it e^-3 X (3^0/0!)

OpenStudy (perl):

yes that would be for P(0)

OpenStudy (anonymous):

okay, thank you...

OpenStudy (anonymous):

this is the 2nd question During the next year, what is the probability of at least 3 months (out of 12) with exactly 1 computer shutdown in each?

OpenStudy (anonymous):

@perl